Unpolarized light is incident on a system of three ideal polarizers. The second polarizer is oriented at an angle of 30.0° with respect to the first, and the third is oriented at an angle of 45.0° with respect to the first. If the light that emerges from the system has an intensity of 31.8 W/m2, what is the intensity of the incident light?
Answer #1

Given:

\theta_{1}=30^{\circ}=\text{angle between the second polarizer and first polarizer}

\theta_{2}=45^{\circ}-30^{\circ}=15^{\circ}=\text{angle between the third polarizer and second polarizer}

I=31.8W/m^{2}=\text{intensity of light coming out from the third polarizer}

I_{o}=?=\text{intensity of unpolarized light}

Applying Malus's Law:

The intensity of light coming out from the third polarizer I=\frac{I_{o}}{2}*\cos^{2}\theta_{1}*\cos^{2}\theta_{2} [where \frac{I_{o}}{2} = intensity of light coming out from the first polarizer].

Therefore, after putting the values in the above equation:

I=\frac{I_{o}}{2}*\cos^{2}\theta_{1}*\cos^{2}\theta_{2}

\Rightarrow 31.8=\frac{I_{o}}{2}*\cos^{2}30^{\circ}*\cos^{2}15^{\circ}

\Rightarrow I_{o}=\frac{2*31.8}{\cos^{2}30^{\circ}*\cos^{2}15^{\circ}}=90.89W/m^{2}

Therefore, the intensity of incident light is 90.89W/m^{2} . [answer]